RC meeting and Run Plan, Friday Oct. 20, 2023

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time: Monday - Friday at 3:30pm

  • Location: Counting house conference room
  • ZoomGov Meeting Connection Information


Useful Links


Hall A call-in instructions:

  • Call RC first.
  • Techs on site 7am-4:30pm Mon-Thu, 7-11am Fri. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
  • After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.

Hall Access procedure:

  • Access check list (binder on the counting room desk) needs to be filled out before making an access.
  • Scan (take pictures with phone) and post it to the logbook.
  • Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.

New instruction for shift crew when changing target:

  • Beam restoration instructions have been updated. Printed version is in the counting room.
  • When we change target, make sure the SOFs of Target BPMs are set to the correct values.
  • SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for major target positions are on the Control Room whiteboard.)

Status Update

Overall Summary


  • Sieve came out and the end of day shift
    • Early analysis of data suggests we have enough stats to be useful for GEM alignment
  • Beam Checkout/Production procedure started early in swing
    • Data showed issue with HCal TDC but was unclear if software or hardware issue
    • Decided to wait until morning and verify that it was not a software issue overnight
    • Escorted access this morning to replace distribution board
    • Seems to look ok now - (https://logbooks.jlab.org/entry/4203728)
  • NMR at 11:30 had weird side peaks on the bottom two plots (https://logbooks.jlab.org/entry/4203757)
    • Had to take a second NMR at this time, looked better - 45.8%
    • 14:30 NMR 48.9 %
  • High power test of the accelerator scheduled for Monday 10/23 from 9 am - 1 pm
    • Beam expected to be very trippy, PD requests we have a plan now if we want glass in the beam during this time or not
    • By this they mean do we want to put the sieve in again or not
  • GEn received another 1 week extension, now runs through 10/31
    • A_LL shifted 1 week, no runtime impacted

Shift Plan

  • Production Data


  • Beam down ~6 hours between 9 pm and 3 am

Opportunistically between spin ups (?):

  • H2 Reference data

Accelerator status

  • Maintenace Day Wednesday 10/25
  • If you want to get into the hall, please communicate with the RC to get this organized!

Data collected :

  • 0.847 C with HCal TDC issue
  • 0.411 C since HCal fix up to 3 pm

Hall A Subsystems Status


  • Laser power and alignment were adjusted Wednesday, pol ~47% in beam
  • Had an NMR issue at 11:30 and had to retake a 2nd one


  • HCal (Jiwan Poudel)
    • TDC Update
  • Bigbite (Arun Tadepalli)
    • Mostly good (Provakar)
    • SC 18-19 missing
  • DAQ (Alex Camsonne)
  • GEMs (Holly Szumila-Vance)
    • Mostly good (Andrew)
    • SBS Layer 2 hole
  • Grinch
    • Looks fine (Eric)
  • Hodoscopes (Rachel Montgomery):
    • Looks fine (Gary)


  • SBS and BB are on.

Hall access requests

If anyone wants to make an access, let the RC know!

Sub Systems

  • SBS magnet and Bigbite magnets
    • work as expected
  • Corrector magnets:
    • work as expected
  • Target status (magnets, measurements, laser, installation, alignment, polarization, power etc)
  • HCal (Jiwan Poudel)
    • TDC Update
  • Bigbite (Arun Tadepalli)
    • Mostly good
    • SC 18-19 missing
  • DAQ (Alex Camsonne)
  • GEMs (Holly Szumila-Vance)
    • Mostly good
    • SBS Layer 2 hole
  • Grinch
    • Looks fine
  • Hodoscopes (Rachel Montgomery):
    • Looks fine

Shift schedule


  • Saturday 10/21 DAY (Target Operator, trade requested)

Please check when your shifts are!


Please sign up for any open shifts!

Running Plan

  • Standard production running at 45uA, with NMRs every 3 hours.

==> List of items in case He3 production running isn’t ready:

  1. Take data on reference cell at 10 atm H2 [85% (4200 elastic events) more data required for Bigbite momentum calibration]
  2. BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on) (? is this different than the optics runs?)